### v1

# Training a Foundation Model for Materials on a Budget

Abstract
Foundation models for materials modeling are advancing quickly, but their training remains expensive, often placing state-of-the-art methods out of reach for many research groups. We introduce Nequix, a compact E(3)-equivariant potential that pairs a simplified NequIP design with modern training practices, including equivariant root-mean-square layer normalization and the Muon optimizer, to retain accuracy while substantially reducing compute requirements. Built in JAX, Nequix has 700K parameters and was trained in 500 A100-GPU hours. On the Matbench-Discovery and MDR Phonon benchmarks, Nequix ranks third overall while requiring less than one quarter of the training cost of most other methods, and it delivers an order-of-magnitude faster inference speed than the current top-ranked model. We release model weights and fully reproducible codebase at https://github.com/atomicarchitects/nequix.
1 Introduction
Machine learned inter-atomic potentials (MLIPs) are rapidly improving in capability and scope, with foundation models trained on broad datasets of atomistic materials offering the promise of augmenting or replacing expensive ab initio density functional theory (DFT) calculations (Batatia et al., 2023). While performance on community benchmarks such as Matbench-Discovery (Riebesell et al., 2025) is rising, the computational costs of both data generation and curation as well as the training of MLIP models on these datasets remain prohibitively expensive for many labs.
We pursue an orthogonal goal to scaling: a lower computational cost recipe that preserves strong downstream accuracy. Concretely, we revisit a simplified E(3)-equivariant architecture based on NequIP (Batzner et al., 2022) with modern training practices: root-mean-square layer normalization for stability, dynamic batching to maximize GPU utilization, and optimizer choices inspired by “speedrunning” deep learning workflows (Jordan et al., 2024a). The resulting model, Nequix, has 700K parameters and can be trained in 500 GPU hours, while remaining competitive with larger and more costly to train models on Matbench-Discovery and other phonon prediction tasks.
[…]
2 Methods
Figure 1: (a) Nequix architecture, a simplified version of NequIP (Batzner et al., 2022), with a species-independent residual connection and layer normalization. (b) Combined performance scores of compliant models on the Matbench-Discovery (unique prototypes subset). (c) Available published training times of current compliant models.
Architecture
Nequix follows a simplified version of the NequIP (Batzner et al., 2022) architecture, as shown in figure 1a. We adopt two modifications suggested by Park et al. (2024): the species-specific self-connection layer within the interaction block is replaced with a single linear layer, and unused non-scalar representations are discarded from the final layer. Lastly, we add an equivariant root-mean-square layer normalization (RMSNorm) Liao et al. (2023), which we find improves performance in our optimization setting. We document the full architecture hyper-parameters and the rationale behind each decision in Table A.1.
Implementation
Nequix is implemented in JAX (Bradbury et al., 2018) with Equinox (Kidger and Garcia, 2021), taking advantage of just-in-time compilation and efficient automatic differentiation. Following standard energy-conserving MLIP practice (Fu et al., 2025), forces are obtained as the negative energy gradient with respect to atomic positions, $\nabla_{\mathbf{r}}E$, and stresses as the energy derivative with respect to strain, normalized by volume, $\sigma=V^{-1}\,\partial E/\partial\varepsilon$, where $E$ is the predicted total energy of the system, $\mathbf{r}$ is an atom position, $\varepsilon$ is the strain tensor, $V$ is the simulation cell volume, and $\sigma$ denotes the stress tensor.
Dynamic batching
Materials radius graphs vary widely in their numbers of nodes and edges due to differences in unit-cell size and atom count. With fixed-size batching, the largest graphs dictate memory usage, leaving GPUs underutilized for most batches. To keep batch workloads more uniform while respecting memory limits, we use dynamic batching (Speckhard et al., 2025): each batch is filled up to caps on the total nodes and edges. We set these caps to $1.1\times$ (ideal batch size) $\times$ (dataset-average nodes-per-graph and edges-per-graph), respectively. Batches are then padded to these caps to fulfill the static shape requirements of JAX (Godwin* et al., 2020).
Optimization and normalization
We compare the widely used Adam optimizer Kingma and Ba (2014) with the recently proposed Muon optimizer Jordan et al. (2024b), which uses the Newton-Schulz algorithm to orthogonalize weight updates. Using a smaller version of Nequix on MPtrj (Jain et al., 2013; Deng et al., 2023) with hidden irreps of 128x0e + 64x1o, we sweep learning rates of $\{0.03,0.01,0.003,0.001\}$ for Adam and Muon, each with and without RMSNorm. The lowest validation error runs for each optimizer are shown in Fig. 2. We find that the Muon configuration achieves comparable energy/force errors to Adam in 60-70% of the epochs, and results in a $7\%$ reduction in energy MAE. We also find a significant reduction in the variance of stress error, which we notice in runs that use RMSNorm. Notably, the presence of the RMSNorm layer generally resulted in lower validation error for Muon-based training configurations, and higher for those using Adam.
Training procedure
The final Nequix model is trained for 100 epochs on MPtrj (Jain et al., 2013; Deng et al., 2023), of which we hold out 5% for validation. More details on the training settings are provided in Sec. A.1. The model was trained on 4 NVIDIA A100 GPUs in 125 hours, for a total cost of 500 GPU hours.
Figure 2: Validation metrics during training of a smaller version of Nequix configuration with Adam and Muon, trying learning rates in $\{0.03,0.01,0.003,0.001\}$ and with/without RMSNorm. This model configuration uses the same hyperparameters as the final model, except with hidden irreps of 128x0e + 64x1o. The dotted horizontal line shows the best validation performance reached during the Adam training.
3 Experiments
3.1 Matbench-Discovery benchmark
Matbench-Discovery (Riebesell et al., 2025) provides a standard framework for evaluating interatomic potentials in a high-throughput materials screening task consisting of geometry optimization and energy prediction on a set of 257,487 generated structures, and thermal conductivity prediction on a set of 103 structures. Ground truth is calculated with DFT/PBE level of theory, the same as MPtrj. The primary metrics include: 1) the F1 for stable/unstable classification after relaxation; 2) root mean squared displacement (RMSD) between predicted and reference structures after relaxation; and 3) symmetric relative mean error in predicted phonon mode contributions to thermal conductivity $\kappa$ ($\kappa_{\mathrm{SRME}}$). A normalized and weighted combination of these metrics are then used to compute a combined performance score (CPS-1), which is used for ranking.
Following Riebesell et al. (2025), we integrate our interatomic potential as Atomic Simulation Environment (ASE) calculator, which is then used to perform structure relaxation and phonon calculations with the default settings of the benchmark. For comparison, we consider only models in the compliant subset of the benchmark. This consists only of models that are trained on MPtrj or subsets, which limits data leakage and offers a more fair comparison among methods. Table 1 contains the performance of Nequix along with all current compliant models at the time of writing. We also include the reported training cost for the models when available, visualized in Fig. 1. We find that Nequix ranks third by CPS-1, outperforming most models at a fraction of the training cost. It is noteworthy that this high ranking is due to high performance in the thermal conductivity task, however, the F1 score is still comparable to many of the other methods.
Table 1: Matbench-Discovery v1 compliant leaderboard, sorted by combined performance score (CPS-1). Metrics are shown for the unique prototypes subset. Train cost is measured in A100 hours. Data as of 2025-08-17.
[…]
eSEN-30M-MP 30.1M - 0.075 0.340 0.831 0.797
Eqnorm MPtrj 1.31M 2000 0.084 0.408 0.786 0.756
Nequix 708K 500 0.085 0.446 0.750 0.729
DPA-3.1-MPtrj 4.81M - 0.080 0.650 0.803 0.718
SevenNet-l3i5 1.17M - 0.085 0.550 0.760 0.714
HIENet 7.51M 2888 0.080 0.642 0.777 0.707
MatRIS v0.5.0 MPtrj 5.83M - 0.077 0.861 0.809 0.681
GRACE-2L-MPtrj 15.3M - 0.090 0.525 0.691 0.681
MACE-MP-0 4.69M 2600 0.092 0.647 0.669 0.644
eqV2 S DeNS 31.2M - 0.076 1.676 0.815 0.522
ORB v2 MPtrj 25.2M - 0.101 1.725 0.765 0.470
M3GNet 228K - 0.112 1.412 0.569 0.428
CHGNet 413K - 0.095 1.717 0.613 0.400
3.2 MDR phonon benchmark
Performance is also evaluated on the MDR phonon benchmark (Loew et al., 2025), a set of 10,000 phonon calculations also done with DFT/PBE level of theory. We follow the identical procedure to Loew et al. (2025), first performing a geometry relaxation, then phonon calculations using displacements of supercells. We report the mean absolute error (MAE) of properties derived from the phonon calculation: maximum phonon frequency $\omega_{\max}$, vibrational entropy $S$, Helmholtz free energy $F$, and heat capacity at constant volume $C_{V}$. Table 2 demonstrates the performance of Nequix compared to other MPtrj-trained models. Similarly to Matbench-Discovery, we achieve performance within the top three of models, with a fraction of the parameter count of other methods.
Table 2: Model performance of MPtrj-trained models on the MDR phonon benchmark, sourced from Loew et al. (2025) and Fu et al. (2025). Metrics are MAE of maximum phonon frequency $\omega_{\max}$ (K), vibrational entropy $S$ (J/K/mol), Helmholtz free energy $F$ (kJ/mol) and heat capacity at constant volume $C_{V}$ (J/K/mol).
Model MAE($\omega_{\max}$) MAE($S$) MAE($F$) MAE($C_{V}$)
eSEN–30M 21 13 5 4
SevenNet–l3i5 26 28 10 5
Nequix 26 33 12 6
SevenNet–0 40 48 19 9
GRACE–2L (r6) 40 25 9 5
MACE 61 60 24 13
CHGNet 89 114 45 21
M3GNet 98 150 56 22
3.3 Inference speed
Figure 3: Inference speed of various models in steps per day.
[…]
4 Conclusion
We presented Nequix, an E(3)-equivariant interatomic potential that pairs a simplified NequIP architecture with modern training practices. Our results show that Nequix achieves competitive accuracy on Matbench-Discovery and the MDR phonon benchmark at less than one quarter of the reported training cost of many contemporaries. This resource-efficient recipe provides a practical alternative to large-scale foundation models and helps broaden access to high-quality atomistic modeling in settings with more limited compute. We release trained weights and a JAX codebase to streamline reuse and extension.
Looking ahead, we see several promising directions: scaling training duration and data while maintaining budget discipline, exploring pretraining and fine-tuning regimes across broader datasets, and pushing cost even lower through model distillation, pruning, quantization, kernel implementations, or more data-efficient training. We hope Nequix serves as a strong, efficient baseline for future work on accessible materials foundation models.
Acknowledgments and Disclosure of Funding
We thank Mit Kotak for helpful discussions and insights. This work was supported by the National Science Foundation under Cooperative Agreement PHY-2019786 (The NSF AI Institute for Artificial Intelligence and Fundamental Interactions, http://iaifi.org/) and by DOE ICDI grant DE-SC0022215. This research used resources of the National Energy Research Scientific Computing Center (NERSC), a Department of Energy User Facility using NERSC award ERCAP0033254.
Appendix A Appendix
A.1 Training and model configuration
Table A.1 shows the hyper-parameters used to train Nequix. The model is trained for 100 epochs, using an MAE loss function on energy and stress, and $l_{2}$ loss on forces. We use a linear warmup with cosine decay learning rate schedule. Figure A.1 shows the energy, force, and stress MAE on the validation set throughout training. The final MAEs are 10.05 meV/atom, 32.79 meV/Å, and 0.22 meV/Å3/atom for energy, forces, and stress respectively.
Table A.1: Hyper-parameters used and rationale behind selection
Hyper-parameter Value Notes/Rationale
Radial cutoff 6 Å Most models use 5 or 6 Å; 6 performed slightly better in preliminary validation performance.
Hidden irreps 128x0e + 64x1o + 32x2e + 32x3o From SevenNet-l3i5.
$L_{\mathrm{max}}$ 3 Consistent with hidden irreps.
$N_{\mathrm{layers}}$ 4 Balance of performance and efficiency.
Radial basis size 8 From NequIP and analysis from Sec. 5.2 of Fu et al. (2025)
Radial MLP size 64 From NequIP.
Radial MLP layers 2 From NequIP.
Polynomial cutoff pp 6.0 From NequIP.
Radial basis function Bessel From NequIP. Also tried Gaussian, which had minimal difference on validation performance.
Learning rate 0.01 Selected from $\{0.03,0.01,0.003,0.001\}$ based on validation performance early in training.
Warmup epochs 0.1 From eSEN.
Warmup factor 0.2 From eSEN.
Optimizer Muon See Sec. 2.
Weight decay 0.001 From eSEN. Also tried 0.0, which led to worse validation performance.
Energy weight 20 From eSEN.
Force weight 20 From eSEN.
Stress weight 5 From eSEN.
Batch size 256 (dynamic) See Sec. 2
Number of epochs 100 Standard training duration.
Figure A.1: Validation curves for Nequix training on MPtrj.

### v2

@@
 We introduce Nequix, a compact E(3)-equivariant potential that pairs a simplified NequIP design with modern training practices, including equivariant root-mean-square layer normalization and the Muon optimizer, to retain accuracy while substantially reducing compute requirements.
-Built in JAX, Nequix has 700K parameters and was trained in 500 A100-GPU hours.
-On the Matbench-Discovery and MDR Phonon benchmarks, Nequix ranks third overall while requiring less than one quarter of the training cost of most other methods, and it delivers an order-of-magnitude faster inference speed than the current top-ranked model.
+Nequix has 700K parameters and was trained in 100 A100 GPU-hours.
+On the Matbench-Discovery and MDR Phonon benchmarks, Nequix ranks third overall while requiring a 20 times lower training cost than most other methods, and it delivers two orders of magnitude faster inference speed than the current top-ranked model.
 We release model weights and fully reproducible codebase at https://github.com/atomicarchitects/nequix.
@@
 We pursue an orthogonal goal to scaling: a lower computational cost recipe that preserves strong downstream accuracy.
-Concretely, we revisit a simplified E(3)-equivariant architecture based on NequIP (Batzner et al., 2022) with modern training practices: root-mean-square layer normalization for stability, dynamic batching to maximize GPU utilization, and optimizer choices inspired by “speedrunning” deep learning workflows (Jordan et al., 2024a).
-The resulting model, Nequix, has 700K parameters and can be trained in 500 GPU hours, while remaining competitive with larger and more costly to train models on Matbench-Discovery and other phonon prediction tasks.
+Concretely, we revisit a simplified E(3)-equivariant architecture based on NequIP (Batzner et al., 2022) with modern training practices: root-mean-square layer normalization for stability, latest custom CUDA kernels (Bharadwaj et al., 2025), and optimizer choices inspired by “speedrunning” deep learning workflows (Jordan et al., 2024a).
+The resulting model, Nequix, has 700K parameters and can be trained in 100 GPU hours, while remaining competitive with larger and more costly to train models on Matbench-Discovery and other phonon prediction tasks.
 […]
@@
 Figure 1: (a) Nequix architecture, a simplified version of NequIP (Batzner et al., 2022), with a species-independent residual connection and layer normalization.
-(b) Combined performance scores of compliant models on the Matbench-Discovery (unique prototypes subset).
+(b) Combined performance scores of compliant models on the Matbench-Discovery (unique prototypes subset), collected on 2025-08-17.
 (c) Available published training times of current compliant models.
@@
 Implementation
-Nequix is implemented in JAX (Bradbury et al., 2018) with Equinox (Kidger and Garcia, 2021), taking advantage of just-in-time compilation and efficient automatic differentiation.
+Nequix is implemented in both JAX (Bradbury et al., 2018; Kidger and Garcia, 2021) and PyTorch(Paszke et al., 2019; Ansel et al., 2024), taking advantage of just-in-time compilation and efficient automatic differentiation.
 Following standard energy-conserving MLIP practice (Fu et al., 2025), forces are obtained as the negative energy gradient with respect to atomic positions, $\nabla_{\mathbf{r}}E$, and stresses as the energy derivative with respect to strain, normalized by volume, $\sigma=V^{-1}\,\partial E/\partial\varepsilon$, where $E$ is the predicted total energy of the system, $\mathbf{r}$ is an atom position, $\varepsilon$ is the strain tensor, $V$ is the simulation cell volume, and $\sigma$ denotes the stress tensor.
@@
 Batches are then padded to these caps to fulfill the static shape requirements of JAX (Godwin* et al., 2020).
+This capability is only used for the JAX implementation.
 Optimization and normalization
@@
 Notably, the presence of the RMSNorm layer generally resulted in lower validation error for Muon-based training configurations, and higher for those using Adam.
+GPU kernels
+There has been recent work (Bharadwaj et al., 2025; NVIDIA, 2024; Tan et al., 2025; Lee et al., 2025) on writing custom GPU kernels for the expensive equivariant tensor product operation (Xie et al., 2025).
+These kernels fuse the tensor product and the outer gather-scatter from the message passing step into a single GPU kernel.
+This avoids storing costly edge-based intermediates in GPU memory, improving both runtime and memory usage.
 Training procedure
@@
 A.1.
-The model was trained on 4 NVIDIA A100 GPUs in 125 hours, for a total cost of 500 GPU hours.
+The model was trained on 2 NVIDIA A100 80 GB GPUs in 50 hours, for a total cost of 100 GPU hours.
 Figure 2: Validation metrics during training of a smaller version of Nequix configuration with Adam and Muon, trying learning rates in $\{0.03,0.01,0.003,0.001\}$ and with/without RMSNorm.
@@
 Eqnorm MPtrj 1.31M 2000 0.084 0.408 0.786 0.756
-Nequix 708K 500 0.085 0.446 0.750 0.729
+Nequix 708K 100 0.085 0.446 0.750 0.729
 DPA-3.1-MPtrj 4.81M - 0.080 0.650 0.803 0.718
@@
 Model MAE($\omega_{\max}$) MAE($S$) MAE($F$) MAE($C_{V}$)
-eSEN–30M 21 13 5 4
-SevenNet–l3i5 26 28 10 5
+eSEN-30M 21 13 5 4
+SevenNet-l3i5 26 28 10 5
 Nequix 26 33 12 6
-SevenNet–0 40 48 19 9
-GRACE–2L (r6) 40 25 9 5
+SevenNet-0 40 48 19 9
+GRACE-2L (r6) 40 25 9 5
 MACE 61 60 24 13
@@
 Figure 3: Inference speed of various models in steps per day.
+Finally, we compare inference speed with existing interatomic potentials by using ASE Larsen et al.
+(2017) to run calculations on diamond across varying unit cell sizes, timing the calculations for each model.
+[^1] See https://github.com/mitkotak/matbench-speed for inference speed benchmarking code..
+We run each model in the default configuration in which it used within its ASE calculator, with compilation and kernels wherever specified in the documentation.
+See Section A.2 for more information on benchmarking setup.
 […]
@@
 This resource-efficient recipe provides a practical alternative to large-scale foundation models and helps broaden access to high-quality atomistic modeling in settings with more limited compute.
-We release trained weights and a JAX codebase to streamline reuse and extension.
+We release trained weights and a JAX/PyTorch codebase to streamline reuse and extension.
 Looking ahead, we see several promising directions: scaling training duration and data while maintaining budget discipline, exploring pretraining and fine-tuning regimes across broader datasets, and pushing cost even lower through model distillation, pruning, quantization, kernel implementations, or more data-efficient training.
@@
 Acknowledgments and Disclosure of Funding
-We thank Mit Kotak for helpful discussions and insights.
-This work was supported by the National Science Foundation under Cooperative Agreement PHY-2019786 (The NSF AI Institute for Artificial Intelligence and Fundamental Interactions, http://iaifi.org/) and by DOE ICDI grant DE-SC0022215.
+This work was supported by the National Science Foundation under Cooperative Agreement PHY-2019786 (The NSF AI Institute for Artificial Intelligence and Fundamental Interactions, http://iaifi.org/) NSF Graduate Research Fellowship program under Grant No.
+DGE-1745302, and by DOE ICDI grant DE-SC0022215.
 This research used resources of the National Energy Research Scientific Computing Center (NERSC), a Department of Energy User Facility using NERSC award ERCAP0033254.
@@
 Figure A.1: Validation curves for Nequix training on MPtrj.
+A.2 Inference Configuration
+We summarize all of the inference configurations in Table A.2.
+We use MPTrj versions for all of the models except NequIP where the MPtrj model did not have kernel support, so we instead use the OMat24 model after confirming that they have the same hyperparams.
+For compatibility with eSEN, we use the older fairchem OCP calculator, and expect the latest version to have less overhead.
+Table A.2: Inference configuration for the ASE benchmarking setup
+Model Compile Kernels
+eSEN-30M $\times$ $\times$
+MACE-MP0 torch.compile cuEquivariance
+Nequix torch.compile openEquivariance